OpenStudy (anonymous):

Solve te system by graphing y= -x + 4 y=3x + 4

OpenStudy (anonymous):

To graph y = -x +4 we have  Can you graph y = (3/1)x + 4

hero (hero):

In general, when solving systems by graphing, you must plot each line separately. Both equations presented are already in the form y = mx + b, where m = slope, and b = the y-intercept. The first line will have a negative slope. The second line has a positive slope. To plot each line follow these steps: 1. First plot the y-intercept 2. To plot the next point, use the the corresponding slope of the line you are plotting. 3. Once you have plotted at least two points, you can draw a straight line through the points 4. Label the line accordingly in the form y = mx + b In this particular case, we're asked to plot y = -x+4 and y = 3x+4 To plot y = -x + 4: 1. Plot point (0,4), the y-intercept 2. Since the slope of the line is -1, start at point (0,4), then move 1 unit to the right, and then one unit downward. Your next point will be (1,3) 3. Draw a line through points (0,4) and (1,3) 4. Label the line you've just created y = -x + 4 To plot y = 3x + 4 1. Plot point (0,4), the y-intercept 2. Since the slope of the line is 3, start at point (0,4), the move one unit to the right, and three units up. Your next point will be (1,7) 3. Draw a line through points (0,4) and (1,7) 4. Label the line you've just created y = 3x+4 Notice that both lines intersect at (0,4). Also recognize in general that when two lines have the same y-intercept, the interception point will be the solution of the system. The graph of the solution should look something like this when done: